a body falling froma vertical height of 10 m pierces through a distance of 1 m in sand.  average retadation in sand faced by body is ..a) g      b) 9g    c) 100g    d) 1000g

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Velocity of body just before hitting the sand=\sqrt{2gh}=10\sqrt{2}m/s

The retarding force exerted by the sand onto the body is opposite to that of acceleration due to gravity.
Hence inside the sand, using equation of motion, one can say,

v^2-u^2=2as
\implies a=-\dfrac{u^2}{2s}=-10g